A baseball is seen to pass upward by a window with a vertical speed v m/s. If the ball was thrown by a person d m below on the street, derive expressions for:

1.) the initial speed of the baseball
2.) the maximum altitude reached by the baseball
3.) the time taken by the baseball to reach the street again
1
Expert's answer
2020-06-12T11:23:29-0400

1)


"v_i=\sqrt{V^2+2gd}"

2)


"H=h+\frac{V^2}{2d}"

3)


"-v_i=v_i-gt"

"t=\frac{2v_i}{g}=\frac{2}{g}\sqrt{V^2+2gd}"